4x+2y=12 into slope intercept form

Found 2 solutions by Alan3354, Theo:
Answer by Alan3354(69443)

Solve for y.

Answer by Theo(13342)

start with 4x + 2y = 12
subtract 4x from both sides of the equation to get 2y = 12 - 4x
divide both sides of the equation by 2 to get y = 6 - 2x
arrange the right side expression by descending order of degree to get y = -2x + 6
that would be the slope intercept form of the equation.
it is equivalent to the standard form of the equation and will graph the same line on a graphing calculator that can work with both forms, such as the desmos.com calculator i used to show you.
$$$